Starting from the species mass balance Pa = -5 (pava) +re at derive the reaction-diffusion equation Pdt =pDag V-wa+r where D is diffusivity.

Identify all variables and state all assumptions
Answer #1

Assumption Si- o fluids @ fluids is incompressible is stationary liquids or solid. Notation : relocity ola 8 mass concentratioso -0.60, )+Yer o we have mass elon wort mass averge velocity ja = Pa Chat Java Sor Par Pada = ja + Pop - @ put € in eqn o +o parte +3] obo) = D. Son @ DE put y in Dla * DAB #DARTZ Patra -6 :. Pa=Waß - 6 put 6 in (5)8 DWO = 8 Dord? Watra 마 assumed .: we solidd or stationary liquide velocity components in derivative are zero substantial 1.I have used DAB instead of D(ALPHA)(BETA),please apologise and replace.I have clearly mentioned the assumption and derived the above equation.